Two angles of a triangle are 33 degrees and 42 degrees. What is the measure of the third?

Answer by Theo(13342)

sum of the angles of a triangle is 180 degrees.
33 + 42 = 75
180 - 75 = 105
the third angle is 105 degrees.
105 + 33 + 42 = 180